OpenStudy (anonymous):

If a 2.5 g sample of gold is hammered into an equilateral triangle .28mm thick, how long are the sides?

OpenStudy (anonymous):

Could you help me? @phi

OpenStudy (anonymous):

I think I know how to do it, but I would like some assistance if possible! @phi

OpenStudy (phi):

what is the density of gold ?

OpenStudy (anonymous):

is it 19.3?

OpenStudy (phi):

I don't know, but you should look it up and post it here. and be sure to add the units

OpenStudy (anonymous):

It is 19.3

OpenStudy (anonymous):

I just looked it up, bur I don't know how it would be that.

OpenStudy (phi):

and what are the units ? any idea ?

OpenStudy (anonymous):

I think the final units must be in cm^3 or g/cm^3

OpenStudy (phi):

ok, so we can write a ratio \[ \frac{19.3 \ g}{1 \ cm^3}= \frac{2.5 \ g }{x} \]

OpenStudy (anonymous):

is this using the factor-label method?

OpenStudy (phi):

can you find the value for x? I would first flip both sides, and write the ratio as \[ \frac{1 \ cm^3}{19.3 \ g}= \frac{x}{2.5 \ g } \]

OpenStudy (phi):

I would then multiply both sides by 2.5 g \[ x = \frac{2.5}{19.3}\ cm^3\]

OpenStudy (anonymous):

Would it be .1295?

OpenStudy (anonymous):

Because that's what I did, but then after I did that I got confused

OpenStudy (phi):

yes, but you should always include the units x= 0.12953 cm^3 you have that many cubic cm of gold now we need the volume of an equilateral shape any idea on the formula for its volume ?

OpenStudy (phi):

I would use area of the base * height what is the area of an equilateral triangle? (there is a formula where all you need is the length of one side)

OpenStudy (anonymous):

I thought that the .1295 was the volume? It isn't?

OpenStudy (phi):

it is the volume of the gold and when you pound it into the shape of a triangle, the triangle will have the same volume